# Deep learning methods for inverse problems using connections between proximal operators and Hamilton–Jacobi equations

### v1

@@ 1 Introduction
Inverse problems are ubiquitous mathematical problems that primarily aim at recovering model parameters from noisy data. They arise in many scientific and engineering applications for, e.g., recovering an image from noisy measurements, deblurring, tomographic reconstruction, and compressive sensing [AEOV2023, bertero2021introduction, isakov2017inverse, arridge2019solving]. Since inverse problems are often ill-posed, it is essential to include regularization or prior information about the underlying model or unknown variables. Proximal operators are central to this: they provide a flexible and computationally convenient way to encode priors and to build efficient iterative algorithms (e.g., proximal (sub)gradients, the alternating direction method of multipliers, and other splitting methods). More recently, proximal operators have become key ingredients for state-of-the-art machine learning methods, e.g., plug-and-play methods that replace explicit regularizers by learned denoisers [hu2023plug, jia2025plug], and deep neural architectures that parameterize proximal maps or their gradients, such as learned proximal networks (LPNs) [fang2024whats]. These developments have made proximal methods practical and powerful computational tools.
@@
The recent work of Gribonval and Nikolova [gribonval2020characterization] in nonsmooth optimization has extended the characterization of proximal operators with convex priors to those with nonconvex priors, showing in particular they are functions that are subdifferentials of certain convex potentials. These properties, in particular, were used in [fang2024whats] to develop new deep learning methods, called learned proximal networks (LPNs), to learn from data the underlying prior of a proximal operator.
The paper  [gribonval2020characterization] did not, however, discuss the well-established, existing connections between proximal operators and Hamilton–Jacobi Partial Differential Equations (HJ PDEs) [darbon2015convex, darbon2016algorithms, darbon2021bayesian, chaudhari2018deep, osher2023hamilton]. To see these connections, consider the following HJ PDE with quadratic Hamiltonian function and whose initial data is the prior $J$:
@@
Moreover, the viscosity solution satisfies the crucial property that $\bm{x}\mapsto\frac{1}{2}\left\|{\bm{x}}\right\|_{2}^{2}-tS(\bm{x},t)$ is convex; that is, when paired with Eq. 4, the function $\text{prox}_{tJ}(\bm{x})$ is obtained from differentiating a convex function. This formally connects proximal operators to HJ PDEs, which we emphasize was previously known and established, and the (stronger) characterization obtained in [gribonval2020characterization] [^1] To the best our knowledge, this characterization result was unknown in the theory of HJ PDEs..
In this paper, we leverage the theory of viscosity solutions of HJ PDEs to develop novel deep learning methods to learn from data the prior function $J$ in Eq. 2. To describe our approach, consider the case when the solution $(\bm{x},t)\mapsto S(\bm{x},t)$ to the HJ PDE Eq. 3 is known. (We will consider the case when only samples of it are known in the next paragraph.) This problem was investigated in [barron1999regularity, claudel2011convex, colombo2020initial, esteve2020inverse, misztela2020initial]. In particular, [esteve2020inverse] showed that when $\bm{x}\mapsto S(\bm{x},t)$ is uniformly Lipschitz continuous and $\bm{x}\mapsto\frac{1}{2}\left\|{\bm{x}}\right\|_{2}^{2}-tS(\bm{x},t)$ is convex, there exists a prior $J$ that can recover $S(\bm{x},t)$ exactly. Moreover, there is a natural candidate for the prior, obtained by reversing the time in the HJ PDE Eq. 3 and using $(\bm{x},t)\mapsto S(\bm{x},t)$ as the terminal condition. The resulting backward viscosity solution yields the prior $J_{\text{BVS}}\colon\mathbb{R}^{n}\to\mathbb{R}$ which admits the representation formula
@@
We focus in this paper on the case when $\bm{x}\mapsto S(\bm{x},t)$ is unknown but have access to some samples $\{\bm{x}_{k},S(\bm{x}_{k},t),\nabla_{\bm{x}}S(\bm{x}_{k},t)\}_{k=1}^{K}$ with $t$ fixed. We propose to learn the prior $\bm{y}\mapsto J_{\text{BVS}}(\bm{y})$ by leveraging the crucial fact that $\bm{y}\mapsto J_{\text{BVS}}(\bm{y})+\frac{1}{2}\left\|{\bm{y}}\right\|_{2}^{2}$ is convex, thus enabling approaches based on deep learning and convex neural networks.
@@
Related works: Hamilton–Jacobi PDEs are important to many scientific and engineering applications arising in e.g., optimal control [Bardi1997Optimal, fleming2006controlled, mceneaney2006max, parkinson2018optimal] and physics [Caratheodory1965CalculusI, Caratheodory1967CalculusII], inverse problems for imaging sciences [darbon2015convex, darbon2019decomposition, darbon2021bayesian, darbon2021connecting, darbon2022hamilton], optimal transport [meng2024primal, onken2021ot], game theory [BARRON1984213, Evans1984Differential, ruthotto2020machine], and machine learning [chen2024leveraging, zou2024leveraging]. Recent works focus on developing specialized solution methods for solving high-dimensional HJ PDEs, using, e.g., representation formulas or deep learning methods. These specialized methods leverage certain properties of HJ PDEs, including stochastic aspects and representation formulas [bardi1998hopf, mceneaney2006max, darbon2016algorithms, darbon2022hamilton], to approximate solutions to HJ PDEs more accurately and efficiently than general-purpose methods. See, e.g., [meng2022sympocnet, darbon2023neural, darbon2021some, darbon2020overcoming, park2025neural] for recent works along these lines and [meng2025recent] for a review of the state-of-the-art numerical methods for HJ PDEs.
@@
Deep learning methods have become popular for computing solutions to high-dimensional PDEs as well as their inverse problems. They are popular because neural networks can be trained on data to approximate high-dimensional, nonlinear functions using efficient optimization algorithms. They have been used to approximate solutions to PDEs without any discretization with numerical grids, and for this reason they can overcome, or at least mitigate, the curse of dimensionality. There is a fairly comprehensive literature on deep learning methods for solving PDEs in general, e.g., see [beck2020overview, cuomo2022scientific, karniadakis2021physics].
@@
Organization of this paper: We present background information on proximal operators, Hamilton–Jacobi equations, and convex neural networks in Section 2. Next, we discuss recent results concerning the inverse problem for Hamilton–Jacobi equations when the solution is available, and how they relate to proximal operators and learning priors in inverse problems, in Section 3. Our main theoretical results are presented in Section 4, where we study the inverse problem for Hamilton–Jacobi equations when only incomplete information is available about its solution. We suggest via arguments from max-plus algebra theory for Hamilton–Jacobi PDEs how to learn from data the solution to a certain Hamilton-Jacobi–Jacobi terminal value problem, which can then be used as an estimate for learning the prior function in a proximal operator. We present in Section 5 some numerical experiments for learning the initial data of certain Hamilton–Jacobi PDEs using convex neural networks and the theory of inverse Hamilton–Jacobi PDEs. Finally, we summarize our results in Section 6.
@@ 2.1 Proximal operators
If $x>t$, the only minimum is $y^{*}=x+t$. Likewise, if x<tx<-t, the only minimum is $y^{*}=x-t$. In either cases, $S(x,t)=-x-\frac{t}{2}$. If $0<x\leqslant t$, there are two local minimums, 0 and $x+t$, but the global minimum is attained at $x+t$ and yields $S(x,t)=-\frac{t}{2}-x$. Likewise, if $t\leqslant x<0$, there are two local minimums, 0 and xtx-t, but the global minimum is attained at xtx-t and yields $S(x,t)=-\frac{t}{2}+x$. Finally, if $x=0$, there are three local minimums, t-t, 0, and $t$. The global minimums are attained at t-t or $t$, yielding S(0,t)=t/2S(0,t)=-t/2. Hence we find
@@ 2.2.1 Characteristic equations
where $\bm{z}(t)=S(\bm{x}(t),t)$ and $\bm{x}(0)=J(\bm{x}(0)$. Here, $t\mapsto\bm{p}(t)$ is constant with $\bm{p}(t)\equiv\bm{p}(0)\in\mathbb{R}^{n}$. The characteristic line that arises from $\bm{x}(0)\in\mathbb{R}^{n}$ is $\bm{x}(t)=\bm{x}(0)+t\bm{p}(0)$, and so $\bm{z}(t)=\bm{z}(0)-\frac{1}{2}\left\|{\bm{p}(0)}\right\|_{2}^{2}$. Taken together, we find
@@ 2.2.2 Viscosity solutions and the Lax–Oleinik formula
satisfy $S_{1}(x,0)=S_{2}(x,0)=J(x)=0$ and both solve the corresponding HJ PDE almost everywhere.
@@
The (unique) viscosity solution has two important properties. First, the function $\bm{x}\mapsto S(\bm{x},t)$ is (1/t)-semiconcave. This is equivalent to requiring the function $\bm{x}\mapsto\psi(\bm{x},t)$ defined in Eq. 6 to be convex, exactly as stipulated in Theorem 2.1. Second, at any point of differentiability of $\bm{x}\mapsto S(\bm{x},t)$, there holds
@@
Let J(x)=|x|J(x)=-|x| in the (one-dimensional) first-order HJ PDE Eq. 3. The function $J$ is uniformly Lipschitz continuous and, as such, the Lax–Oleinik formula $S(x,t)=-\frac{t}{2}-|x|$ is the unique viscosity solution of the corresponding HJ PDE. Note $\bm{x}\mapsto S(\bm{x},t)$ is differentiable everywhere except at $\bm{x}=0$ and $\text{prox}_{tJ}(\bm{x})=\bm{x}-t\nabla_{\bm{x}}S(\bm{x},t)$ everywhere except at $\bm{x}=0$ (see (7)).
@@ 3.1 Reachability and inverse problems for Hamilton–Jacobi equations
We consider here the inverse problem associated to the HJ initial value problem Eq. 3: given $t>0$ and a function $(\bm{x},t)\mapsto S(\bm{x},t)$, identify the set of initial data $J\colon\mathbb{R}^{n}\ \to\mathbb{R}$ such that the viscosity solution of Eq. 3 coincide with $S(\bm{x},t)$. That is, we wish to characterize the set
@@
Now, assume $(\bm{x},t)\mapsto S(\bm{x},t)$ is reachable. What can be said about the nonempty set $I_{t}(S)$? Since $(\bm{x},t)\mapsto S(\bm{x},t)$ is obtained from evolving forward in time the prior function $J$ from 0 to $t$ according to Eq. 3, a natural approach is to do the opposite: evolve backward in time from $t$ to 0 the function $\bm{x}\mapsto S(\bm{x},t)$. That is, we consider the terminal value problem
@@
Under appropriate conditions, the terminal-value problem Eq. 14 has a unique viscosity solution:
@@
Suppose $\bm{x}\mapsto S(\bm{x},t)$ is uniformly Lipschitz continuous and semiconcave. Then the viscosity solution of the terminal-value problem Eq. 14 exists, is unique, and is given by the representation formula
@@
The viscosity solution of (14) is sometimes called the backward viscosity solution (BVS) to distinguish it from the viscosity solution of the initial value problem Eq. 3. The BVS at $\tau=0$ corresponds to fully evolving backward in time the function $\bm{x}\mapsto S(\bm{x},t)$. In what follows, we write $J_{\text{BVS}}\coloneqq\bm{w}(\cdot,0)$. We can use Eq. 6 to write
@@
Theorem 3.5 stipulates that $J_{\text{BVS}}$ is equal everywhere to $J$ on the set $\mathcal{X}_{t}(S)$ and bounds it from below elsewhere. This is a fundamental consequence of the semiconcavity of $\bm{x}\mapsto S(\bm{x},t)$, which regularizes the backward viscosity solution of Eq. 14. We illustrate this below with the negative absolute value prior.
@@
Let J(x)=|x|J(x)=-|x| in the (one-dimensional) first-order HJ PDE Eq. 3. Recall that the unique viscosity solution is given by the Lax–Oleinik formula $S(x,t)=-\frac{t}{2}-|x|$. We now would like to compute the corresponding unique backward viscosity solution to the terminal-value problem Eq. 14. The solution is well-defined because $\bm{x}\mapsto S(\bm{x},t)$ is uniformly Lipschitz continuous and concave. We have
@@
$\operatorname*{arg\,min}_{y\in\mathbb{R}}\left\{\frac{1}{2t}(x-y)^{2}+|y|\right\}=\begin{cases}x-t,\,&\text{if $x>t$},\\ 0,\,&\text{if $x\in[-t,t]$},\\ x+t,\,&\text{if $x<-t$}.\end{cases}$
@@
$J_{\text{BVS}}(x)=\begin{cases}-x,\,&\text{if $x>t$},\\ -\frac{t}{2}-\frac{x^{2}}{2t},\,&\text{if $x\in[-t,t]$},\\ x,\,&\text{if $x<-t$}.\end{cases}$
@@
$tJ_{\text{BVS}}(x)+\frac{1}{2}x^{2}=\begin{cases}\frac{1}{2}(x-t)^{2}-\frac{t^{2}}{2},\,&\text{if $x>t$},\\ -\frac{t^{2}}{2},\,&\text{if $x\in[-t,t]$},\\ \frac{1}{2}(x+t)^{2}-\frac{t^{2}}{2},\,&\text{if $x<-t$},\end{cases}$
@@
The results here apply when the function $\bm{x}\mapsto S(\bm{x},t)$ is known. What happens when only a finite set of values of this function are available?

### v2

@@ 1 Introduction
Inverse problems are ubiquitous mathematical problems that primarily aim at recovering model parameters from noisy data. They arise in many scientific and engineering applications for, e.g., recovering an image from noisy measurements, deblurring, tomographic reconstruction, and compressive sensing [AEOV2023, arridge2019solving, bertero2021introduction, isakov2017inverse]. Since inverse problems are often ill-posed, it is essential to include regularization or prior information about the underlying model or unknown variables. Proximal operators are central to this: they provide a flexible and computationally convenient way to encode priors and to build efficient iterative algorithms (e.g., proximal (sub)gradients and splitting methods). More recently, proximal operators have become key ingredients for state-of-the-art machine learning methods, such as plug-and-play methods that replace explicit regularizers by learned denoisers [hu2023plug, jia2025plug] or deep neural architectures parameterizing proximal maps or their gradients. These developments have made proximal methods practical and powerful computational tools.
@@
The recent work of Gribonval and Nikolova [gribonval2020characterization] in nonsmooth optimization has extended the characterization of proximal operators with convex priors to those with nonconvex priors, showing they are functions that are subdifferentials of certain convex potentials. These properties, in particular, were used in [fang2024whats] to develop new deep learning methods, called learned proximal networks (LPNs), to learn from data the underlying prior of a proximal operator.
The paper [gribonval2020characterization] did not, however, discuss the well-established, existing connections between proximal operators and Hamilton–Jacobi Partial Differential Equations (HJ PDEs) [chaudhari2018deep, darbon2015convex, darbon2021bayesian, darbon2016algorithms, osher2023hamilton]. To see these connections, consider the following HJ PDE with quadratic Hamiltonian function and whose initial data is the prior $J$:
@@
Finally, the viscosity solution is semiconcave, that it, the function $\bm{x}\mapsto\frac{1}{2}\left\|{\bm{x}}\right\|_{2}^{2}-tS(\bm{x},t)$ is convex. Thus, when paired with Eq. 4, the function $\text{prox}_{tJ}(\bm{x})$ is obtained from differentiating a convex function. This formally connects proximal operators to HJ PDEs (which we emphasize was previously known and established) and the stronger [^1] To the best our knowledge, this characterization result was unknown in the theory of HJ PDEs. characterization obtained in [gribonval2020characterization].
In this paper, we propose to leverage the theory of viscosity solutions of HJ PDEs to develop novel deep learning methods for learning the prior function $J$ in Eq. 2 directly from data. To describe our approach, suppose the solution $(\bm{x},t)\mapsto S(\bm{x},t)$ to the HJ PDE Eq. 3 is known. (We will consider the case when only samples of it are known later.) This problem was investigated in [barron1999regularity, claudel2011convex, colombo2020initial, esteve2020inverse, misztela2020initial]. In particular, [esteve2020inverse] showed that when $\bm{x}\mapsto S(\bm{x},t)$ is uniformly Lipschitz continuous and $\bm{x}\mapsto\frac{1}{2}\left\|{\bm{x}}\right\|_{2}^{2}-tS(\bm{x},t)$ is convex, there exists a prior $J$ recovering $S(\bm{x},t)$ exactly. Moreover, there is a natural choice for the prior, obtained by reversing the time in the HJ PDE Eq. 3 and using $(\bm{x},t)\mapsto S(\bm{x},t)$ as the terminal condition. The resulting backward viscosity solution yields the prior $J_{\text{BVS}}\colon\mathbb{R}^{n}\to\mathbb{R}$ which admits the representation formula
@@
We shall focus in this paper on the case where $\bm{x}\mapsto S(\bm{x},t)$ is unknown but have access to some samples $\{\bm{x}_{k},S(\bm{x}_{k},t),\nabla_{\bm{x}}S(\bm{x}_{k},t)\}_{k=1}^{K}$ with $t$ fixed. We propose to learn the prior $\bm{y}\mapsto J_{\text{BVS}}(\bm{y})$ by leveraging the crucial fact that $\bm{y}\mapsto J_{\text{BVS}}(\bm{y})+\frac{1}{2}\left\|{\bm{y}}\right\|_{2}^{2}$ is convex, thus enabling approaches based on deep learning and convex neural networks.
@@
Related works: Hamilton–Jacobi PDEs are important to many scientific and engineering applications arising in e.g., optimal control [Bardi1997Optimal, fleming2006controlled, mceneaney2006max, parkinson2018optimal] and physics [Caratheodory1965CalculusI, Caratheodory1967CalculusII], inverse problems for imaging sciences [darbon2015convex, darbon2021bayesian, darbon2019decomposition, darbon2022hamilton], optimal transport [meng2024primal, onken2021ot], game theory [BARRON1984213, Evans1984Differential, ruthotto2020machine], and machine learning [chen2024leveraging, zou2024leveraging]. Recent works focus on developing specialized methods for solving high-dimensional HJ PDEs via e.g., representation formulas or deep learning. These specialized methods leverage certain properties of HJ PDEs, including stochastic aspects and representation formulas [bardi1998hopf, darbon2022hamilton, darbon2016algorithms, mceneaney2006max], to approximate solutions to HJ PDEs more accurately and efficiently than general-purpose methods. See, e.g., [darbon2023neural, darbon2020overcoming, darbon2021some, meng2022sympocnet, park2025neural] for recent works along these lines and [meng2025recent] for a review of the state-of-the-art numerical methods for HJ PDEs.
@@
Deep learning methods have become popular for computing solutions to high-dimensional PDEs and their inverse problems. This is because neural networks can be trained on data to approximate high-dimensional, nonlinear functions using efficient optimization algorithms. They have been used to approximate solutions to PDEs without any discretization with numerical grids, and for this reason they can overcome, or at least mitigate, the curse of dimensionality. There is a fairly comprehensive literature on deep learning methods for solving PDEs in general, e.g., see [beck2020overview, cuomo2022scientific, karniadakis2021physics].
@@
Organization of this paper: We present background information on proximal operators, Hamilton–Jacobi equations, and convex neural networks in Section 2. Next, we discuss recent results concerning the inverse problem for Hamilton–Jacobi equations when the solution is available, and how they relate to proximal operators and learning priors in inverse problems, in Section 3. We present some theoretical results in Section 4, where we study the inverse problem for Hamilton–Jacobi equations when only incomplete information is available about its solution. We suggest via arguments from max-plus algebra theory for Hamilton–Jacobi PDEs how to learn from data the solution to a certain Hamilton-Jacobi–Jacobi terminal value problem, which can then be used as an estimate for learning the prior function in a proximal operator. We present in Section 5 some numerical experiments for learning the initial data of certain Hamilton–Jacobi PDEs using convex neural networks and the theory of inverse Hamilton–Jacobi PDEs. Finally, we summarize our results in Section 6.
@@ 2.1 Proximal operators
A short calculation shows that
@@ 2.2.1 Characteristic equations
where $\bm{z}(t)=S(\bm{x}(t),t)$ and $\bm{x}(0)=J(\bm{x}(0))$. Here, $t\mapsto\bm{p}(t)$ is constant with $\bm{p}(t)\equiv\bm{p}(0)\in\mathbb{R}^{n}$. The characteristic line that arises from $\bm{x}(0)\in\mathbb{R}^{n}$ is $\bm{x}(t)=\bm{x}(0)+t\bm{p}(0)$, and so $\bm{z}(t)=\bm{z}(0)-\frac{1}{2}\left\|{\bm{p}(0)}\right\|_{2}^{2}$. Taken together, we find
@@ 2.2.2 Viscosity solutions and the Lax–Oleinik formula
satisfy $S_{1}(x,0)=S_{2}(x,0)=0$ and solve the corresponding HJ PDE.
@@
The unique viscosity solution has two important properties. First, the function $\bm{x}\mapsto S(\bm{x},t)$ is (1/t)-semiconcave, that is, the function $\bm{x}\mapsto\psi(\bm{x},t)$ defined in Eq. 6 is convex, as per Theorem 2.1. Second, at any point of differentiability of $\bm{x}\mapsto S(\bm{x},t)$,
@@
Let J(x)=|x|J(x)=-|x| in the HJ PDE Eq. 3. The function $J$ is uniformly Lipschitz continuous and, as such, the Lax–Oleinik formula $S(x,t)=-\frac{t}{2}-|x|$ is the unique viscosity solution of the corresponding HJ PDE. Note $\bm{x}\mapsto S(\bm{x},t)$ is differentiable everywhere except at $\bm{x}=0$ and $\text{prox}_{tJ}(\bm{x})=\bm{x}-t\nabla_{\bm{x}}S(\bm{x},t)$ everywhere except at $\bm{x}=0$ (see (7)).
@@ 3.1 Reachability and inverse problems for Hamilton–Jacobi equations
We consider the inverse problem associated to the HJ initial value problem Eq. 3: given $t>0$ and a function $(\bm{x},t)\mapsto S(\bm{x},t)$, find the set of initial data $J\colon\mathbb{R}^{n}\ \to\mathbb{R}$ such that the viscosity solution of Eq. 3 is $S(\bm{x},t)$. That is, we wish to find the set
@@
Now, assume $(\bm{x},t)\mapsto S(\bm{x},t)$ is reachable. What can we say about $I_{t}(S)$? Since we evolve forward in time $J$ from 0 to $t$ via Eq. 3 to obtain $S(\bm{x},t)$, a natural approach then is to do the opposite: evolve backward in time $S(\bm{x},t)$ from $t$ to 0 via an appropriate HJ PDE. That is, we consider the terminal value problem
@@
Under certain conditions, the HJ PDE Eq. 14 has a unique viscosity solution:
@@
Suppose $\bm{x}\mapsto S(\bm{x},t)$ is uniformly Lipschitz continuous and semiconcave. Then the HJ PDE Eq. 14 has a unique viscosity solution given by
@@
The viscosity solution of (14) is sometimes called the backward viscosity solution (BVS) to distinguish it from its forward counterpart given by Eq. 3. The BVS at $\tau=0$ corresponds to fully evolving backward in time the function $\bm{x}\mapsto S(\bm{x},t)$. In what follows, we write $J_{\text{BVS}}\coloneqq\bm{w}(\cdot,0)$. We can use Eq. 6 to write
@@
Theorem 3.5 is a fundamental consequence of the semiconcavity of $\bm{x}\mapsto S(\bm{x},t)$, which regularizes the backward viscosity solution of Eq. 14. We illustrate this below with the negative absolute value prior.
@@
Let J(x)=|x|J(x)=-|x| in the HJ PDE Eq. 3. Recall that its unique viscosity solution is given by the Lax–Oleinik formula $S(x,t)=-\frac{t}{2}-|x|$. We now compute the corresponding unique backward viscosity solution to the terminal-value problem Eq. 14. The solution is well-defined because $\bm{x}\mapsto S(\bm{x},t)$ is uniformly Lipschitz continuous and concave. We have
@@
(17) $\operatorname*{arg\,min}_{y\in\mathbb{R}}\left\{\frac{1}{2t}(x-y)^{2}+|y|\right\}=\begin{cases}x-t,\,&\text{if $x>t$},\\ 0,\,&\text{if $x\in[-t,t]$},\\ x+t,\,&\text{if $x<-t$}.\end{cases}$
@@
$J_{\text{BVS}}(x)=\begin{cases}-|x|,\,&\text{if $x<-t$ or $x>t$},\\ -\frac{t}{2}-\frac{x^{2}}{2t},\,&\text{if $x\in[-t,t]$}.\end{cases}$
@@
(18) $tJ_{\text{BVS}}(x)+\frac{1}{2}x^{2}=\begin{cases}\frac{1}{2}(x-t)^{2}-\frac{t^{2}}{2},\,&\text{if $x>t$},\\ -\frac{t^{2}}{2},\,&\text{if $x\in[-t,t]$},\\ \frac{1}{2}(x+t)^{2}-\frac{t^{2}}{2},\,&\text{if $x<-t$},\end{cases}$
@@
The results here apply when the function $\bm{x}\mapsto S(\bm{x},t)$ is known. Can anything be said when only a finite set of values of $\bm{x}\mapsto S(\bm{x},t)$ is available?